# Efficient Learning for Linear Properties of Bounded-Gate Quantum Circuits

### v1

@@ H.2 Proof of Corollary 1
SM J More numerical simulations
In this section, we demonstrate more simulation details omitted in the main text. All numerical simulations in this work are conducted using Python, primarily leveraging the Pennylane and PastaQ libraries. For small qubit sizes with $N\leq 20$, we use Pennylane to generate training examples and make predictions for new inputs. For larger qubit sizes with $N>20$, we utilize PastaQ to generate training data and Numpy to handle post-processing and new input predictions.
@@
The contents in the rest of this section are organized as follows. In SM J.1, we provide more simulation details about predicting properties of rotational $N$-qubit GHZ states. Then, in SM J.2, we illustrate more simulation results about the global Hamiltonian simulation task. Last, in SM J.3, we demonstrate how to use the proposed ML model to enhance variational quantum algorithms, including variational quantum Eigen-solver and quantum neural networks.
J.1 Numerical simulations of $N$-qubit rotational GHZ states
@@
|$\ket{\text{GHZ}(\bm{x})}=\left(\operatorname{RY}_{1}(\bm{x}_{1})\otimes\operatorname{RY}_{N/2}(\bm{x}_{2})\otimes\operatorname{RY}_{N}(\bm{x}_{3})\right)\frac{\ket{0\cdots 0}+\ket{1\cdots 1}}{\sqrt{2}}.$ (143)
The circuit implementation of this class of quantum states is visualized in Fig. J.7(a).
Figure J.7: More simulation results of rotational N-qubit GHZ states. (a) Circuit implementation of the rotational $N$-qubit GHZ states. The three $\operatorname{RY}$ gates apply to the first qubit, the $(N/2)$-th qubit, and the last qubit. (b) Prediction error. The root mean squared (RMS) error of the trained ML model with varied truncation $\Lambda$, the shot number $T$, and the qubit count $N$.
@@
Prediction error versus varied number of qubits $N$. We append more simulation results about how the prediction error of the proposed ML model depends on the shot number $T$, the truncation $\Lambda$, and the qubit counts $N$. To be specific, we fix the number of training examples to be $n=500$ and collect these training examples under different qubit count $N$, where the maximum shot number is set as $T=150$. Fig. J.7(b) visualizes the root mean squared (RMS) prediction error under different settings on $10$ test examples. The achieved results indicate that the performance of the proposed ML model is dominated by the truncation number $\Lambda$ (i.e., the dimension of classical inputs $d$) and not sensitive to the number of qubits, which echoes our theoretical analysis. More precisely, when $\Lambda\geq 2$ and $T\geq 100$, the prediction error attains a very low value for both $N=10,30,60$.
J.2 More details of synthetic global Hamiltonian simulation
Hyper-parameter settings. In the main text, we adopt three settings of the maximum frequency length, i.e., $\Lambda=1,2,3,4$, to evaluate the performance of the proposed ML model. The corresponding cardinality of the frequency set is $|\mathfrak{C}(\Lambda=1)|=61$, $\mathfrak{C}|(\Lambda=2)|=1801$, $|\mathfrak{C}(\Lambda=3)|=472761$, and respectively. When evaluating the statistical performance of the proposed ML model with $n<50000$, we sample a subset from the whole dataset using different random seeds. The random seeds for all relevant simulations are set as $12345$, $22222$, $33333$, $44444$, and $55555$, respectively.
Standard error for the simulation results in Fig. 3(b). For ease of visualization, Fig. 3(b) only demonstrates the averaged RMS prediction error of our model. For completeness, here we append the corresponding standard error in Table 2. Particularly, for all settings with $\Lambda\in\{1,2,3,4\}$, the averaged RMS prediction error and the corresponding standard error continuously decrease with the increased number of training examples.
Table 2: Simulation results of predicting the magnetization of states evolved by the 60-qubit global Hamiltonian.
@@
The role of shot number $T$. Here we conduct numerical simulations to explore how the shot number $T$ influences the prediction error. Specifically, we fix the number of training examples as $n=500$ and the truncation value as $\Lambda=1$, but vary the shot number from $T=50$ to $T=00$. Each setting is repeated five times to collect the statistical results, where the randomness stems from distilling different training datasets from $600$ training examples. The simulation results are summarized in Table. 3, where the averaged prediction error does not decrease with the increased number of measurements. This phenomenon echoes our theoretical analysis, suggesting that once the shot number exceeds a certain threshold, it does not heavily affect the performance of the proposed model.
Table 3: Simulation results of predicting the magnetization of states evolved by the 60-qubit global Hamiltonian with varied shot numbers.
@@
$n=500$ $\Lambda=1$ $0.0130\pm 0.0109$ $0.0251\pm 0.0038$
J.3 More simulation results on enhanced variational quantum algorithms
@@
Algorithmic implementation. Recall that contemporary quantum devices encounter limitations such as connectivity, gate fidelities, and coherence time. To overcome these constraints, experimentalists typically adopt hardware-efficient ansatzes (HEAs) constructed from the native gate set from the specified quantum device to implement various variational quantum algorithms [117, 118]. Essentially, these HEAs are tailored to individual quantum devices, maintaining consistent gate layouts while adjusting parameters for various computational tasks, as delineated by the framework formulated in Eq. (3) of the main text. This intrinsic relation warrants the use of the proposed ML model to enhance plenty of variational quantum algorithms associated with device-specific HEAs.
The learning framework is summarized in Alg. 1. For a given quantum device and its corresponding HEA, the learner first constructs the training dataset following the outlined procedure in the main text. Once the dataset is prepared, the proposed ML model serves as a surrogate to optimize variational quantum algorithms in which the predictions correspond to the gradient information, obviating the need for direct implementation on the specified quantum device.
1. (Dataset construction) Randomly generate classical inputs $\bm{x}\sim[-\pi,\pi]^{d}$, feed it to the specified $N$-qubit quantum device to obtain the pre-measured state $U(\bm{x})\ket{0}^{\otimes N}$, and apply Pauli-based classical shadow with $T$ shots to constitute a single training example $(\bm{x},\tilde{\rho}_{T}(\bm{x}))$;
@@
Since the optimization process (i.e., the third step) is conducted entirely on classical processors, it significantly reduces the overhead associated with accessing sparse quantum devices in the contemporary era.
In what follows, we demonstrate how our approach supports two crucial classes of variational quantum algorithms: the variational quantum Eigen-solvers (VQEs) and quantum neural networks (QNNs). Note that the learning scheme presented below can be readily extended to wide applications covered by variational quantum algorithms.
@@
To estimate the ground state energy $\mathfrak{E}^{*}$, VQE adopts an ansatz $W(\bm{\theta})$ to prepare a variational quantum state |$\ket{\psi(\bm{\theta})}=W(\bm{\theta})\ket{\psi_{0}}^{\otimes N}$ with a fixed input state |$\ket{\psi_{0}}$. The trainable parameters $\bm{\theta}$ are optimized by minimizing the loss function $\mathcal{L}(\bm{\theta},\mathfrak{H})=\operatorname{Tr}\left((\ket{\psi_{0}}\bra{\psi_{0}})W(\bm{\theta})^{\dagger}\mathfrak{H}W(\bm{\theta})\right)$. The optimization of VQE follows an iterative manner, i.e., the classical optimizer continuously leverages the output of the quantum circuits to update $\bm{\theta}$ and the update rule is
@@
$\frac{\partial\mathcal{L}(\bm{\theta}^{(t)},\mathfrak{H})}{\partial\bm{\theta}_{k}}=\frac{1}{2\sin\alpha}\left[\operatorname{Tr}\left((\ket{\psi_{0}}\bra{\psi_{0}})W(\bm{\theta}^{(t,+)})\mathfrak{H}W(\bm{\theta}^{(t,+)})^{\dagger}\right)-\operatorname{Tr}\left((\ket{\psi_{0}}\bra{\psi_{0}})W(\bm{\theta}^{(t,-)})\mathfrak{H}W(\bm{\theta}^{(t,-)})^{\dagger}\right)\right],$ (147)

### v2

@@ H.2 Proof of Corollary 1
SM J Connection and differnces with Ref. [30] and Pauli path simulators
In this section, we detail how our work relates to and differs from Ref. [30] and Pauli path simulators. Specifically, in Subsection J.1, we discuss the complementary relationship between our work and Ref. [30], highlighting how the combination of these studies enriches quantum learning theory and deepens our understanding of the learnability of bounded-gate quantum circuits. In SM J.2, we clarify the distinctions between the proposed learning model and Pauli path simulators, addressing differences at both the technical and application levels.
J.1 Complementation with Ref. [30]
For comprehensive, we summarize the theoretical results achieved by Ref. [30] and our work in Table 2. Particularly, Ref. [30] investigates the computational complexities of quantum learners across three learning tasks:
  • Task (i) learning bounded-gate unitaries;
  • Task (ii) learning bounded-gate quantum states;
  • Task (iii) learning $\mathbb{R}$-valued classical functions with 1-Lipschitz continuity.
Different from Ref. [30], we explore the computational complexities of classical learners for a new task:
  • Task (iv) predicting the expectation values of bounded-gate quantum circuit for a given observable.
The achieved theoretical results under different learning tasks (i.e., Tasks (i)-(iv)) and different learning paradigms (i.e., quantum learners versus classical learners) are complementary, providing the following two crucial insights in the context of the learnability of bounded-gate quantum circuits.
[…]
Before moving on to discuss the two insights, let us emphasize the necessity of exploring both classical and quantum learners. Recall that Ref. [30] focuses on the quantum learner, which may implement the learning model on quantum computers (e.g., quantum neural networks) and require quantum resources during both the training and inference stages. In contrast, our work centers on the classical learner, which implements the learning model on classical computers, requiring quantum resources only during the data collection stage. This means that both the training and inference stages for classical learners are conducted entirely on classical hardware. A comprehensive analysis about the computational complexities of both quantum and classical learners can guide various practical applications and indicate the utility of quantum computers. On the one side, given the foreseeable scarcity of quantum resources, if classical learners can achieve comparable performance to quantum learners for a specific task, they offer undeniable benefits in terms of resource efficiency. On the other side, if quantum learners demonstrate provable computational advantages over classical learners for a given task, this would establish the practical utility of quantum computing for that particular application. For example, Ref. [72] provides a compelling example of how computational capabilities differ between classical and quantum learners, demonstrating that quantum learners can achieve exponential runtime speedups in predicting the properties of ground states.
J.1.1 Insight I: Computational complexity analysis
Based on the achieved computational complexities presented in Table 2, we draw the following key observations in the context of learning bounded-gate quantum circuits.
Observation I. Let us revisit the achieved sample complexity in Tasks (i), (ii), and (iv) summarized in Table 2. That is, despite the decreasing levels of information gain—from learning an unknown unitary, to an unknown state, and finally to unknown expectation values, we observe a consistent scaling behavior for both classical and quantum learners.
[…]
We would like to note that although Ref. [30] and our work use different descriptions of bounded-gate circuits, the results achieved in our study can be directly extended to their framework. This is because our description of bounded-gate circuits can be efficiently transformed into their setting with only a constant overhead. More precisely, a bounded-gate quantum circuit composed of $\hat{G}$ two-qubit gates can be implemented using at most $d=24\hat{G}$ $\operatorname{RZ}$ gates and $G-d=35\hat{G}$ Clifford gates to implement this circuit. An intuition is provided in Fig. 4. Ref. [124] has proved that three CNOT gates, associated with eight single-qubit gates, are necessary and sufficient in order to implement an arbitrary unitary transformation of two qubits. Meanwhile, any single qubit gate can be realized by three $\operatorname{RZ}$ gates and four single-qubit Clifford gates [125]. Taken together, $3\times 8=24$ $\operatorname{RZ}$ gates and $4\times 8+3=35$ Clifford gates can implement an arbitrary unitary transformation of two qubits. Since there are $\hat{G}$ two-qubit gates, $d=24\hat{G}$ $\operatorname{RZ}$ gates and $G-d=35\hat{G}$ Clifford gates are sufficient to implement this circuit.
Supplementary Fig. 4: Decomposition of an arbitrary two-qubit gate into $\operatorname{RZ}$ and Clifford gates. The dashed box highlights the implementation of the single-qubit gates $u_{*}$ and $v_{*}$ with $\in\{1,2,3,4\}$ by $\operatorname{RZ}$ gates and single-qubit Clifford gates.
According to the result of Theorem 1 in our work, when a classical learner is applied to predict the expectation value of a bounded-gate quantum circuit composed of $\hat{G}$ two-qubit gates, $\tilde{O}((B^{2}\hat{G}+B^{2}N\hat{G})/\epsilon)$ and $\tilde{\Omega}((1-\epsilon)\hat{G}/(\epsilon T))$ training examples are sufficient and necessary to achieve a bounded prediction error. This result is consistent with the findings for Tasks (i) and (ii) achieved in Ref. [30], both of which display a linear-in-$\hat{G}$ sample complexity.
The linear scaling behavior with respect to the number of tunable gates across different learners (i.e., quantum versus classical) and tasks (i.e., Tasks (i), (ii), and (iv)) suggests a consistent linear-in-$d$ (or linear-in-$\hat{G}$) sample complexity for classical learners in Tasks (i)&(ii) and for quantum learners in Task (iv). This conjecture is supported by two key facts. First, covering and packing nets, as well as (quantum) hypothesis selection, are robust tools for analyzing the sample complexity of learning bounded-gate quantum circuits, often providing tight bounds [83, 120, 119, 126, 117]. Second, the cardinality of these nets is intrinsically dependent on the number of gates ($d$, $G$, or $\hat{G}$), which leads to similar results across tasks, despite variations in the intended information gain.
Observation II. We next turn to rethink the achieved runtime complexities in Tasks (i), (ii), and (iv) summarized in Table 2. An immediate observation is that, for all tasks, there are no efficient quantum learners for learning bounded-gate unitaries or quantum states [30, Theorems 2&6]. In addition, there are no efficient classical learners capable of accurately predicting the mean-value space of arbitrary bounded-gate quantum circuits (Theorem 1 in the main text). These results are complementary and offer valuable insights into the unresolved issues outlined in Table 2.
First, the inherent limitations of quantum learners in learning bounded-gate unitaries or states strongly suggest that there are no efficient classical learners capable of accomplishing Tasks (i) and (ii). This is because from an information-theoretic perspective, classical learners cannot extract much more information than quantum learners. Consequently, classical learners are unlikely to outperform quantum learners in these tasks.
Second, while neither quantum nor classical learners can efficiently learn arbitrary bounded-gate quantum circuits, it is natural to ask whether specific conditions on the explored bounded-gate circuits could lead to the computational separations, i.e., quantum learners can provide efficient solutions, whereas classical learners would still face computational hardness. Answering this question can enhance our understanding of the learnability of bounded-gate quantum circuits and help identify the practical utility of quantum learners.
J.1.2 Insight II: Learning algorithm design
The computational hardness of learning bounded-gate quantum circuits in Tasks (i), (ii) and (iv) highlights the need to develop advanced learning algorithms with provable computational efficiency under practical and realistic conditions. Both Ref. [30] and our work make initial strides in this direction. Below, we summarize the advancements of each method and outline potential directions for future research.
For Tasks (i) and (ii), Ref. [30] proposes efficient learning algorithms for relatively restrictive scenarios where the number of two-qubit gates $\hat{G}$ scales logarithmically with the number of qubits $N$, i.e., $\hat{G}\sim\mathcal{O}(\log(N))$. Conceptually, at the inference stage, given any new quantum circuit composed of $\mathcal{O}(\log(N))$ two-qubit gates, the proposed algorithm first identifies the non-trivial $\mathcal{O}(\log(N))$ qubits on which two-qubit gates are applied and then employs Pauli-based tomography on these qubits. The achieved result naturally raises two future research directions. The first direction is to explore whether there exist efficient learning algorithms beyond the logarithmic scaling scenario. The second research direction is to explore how to design effective classical learners that achieve performance comparable to quantum learners in Tasks (i) and (ii).
For Task (iv), we devise an efficient classical-shadow predictor based on the kernel methods. While a direct comparison may not be entirely fair, it is worth highlighting that the conditions required for provable efficiency in our approach are much less restrictive than those in Ref. [30]. Specifically, our protocol can efficiently learn a broad class of quantum circuits composed of a polynomial number of gates, as opposed to being limited to logarithmic scenarios. The achieved results, combined with the unsolved issues in Table 2, lead to two interesting questions to be further explored. The first one is to explore alternative efficient classical learning algorithms. The second one is to analyze the computational separation between quantum and classical learning algorithms for Task (iv).
In summary, the complementary roles of Ref. [30] and our work expand the landscape of learning algorithms for bounded-gate quantum circuits, paving the way for deeper exploration of their utility and limitations in quantum computing.
J.2 Connections and differences with Pauli path simulators
The primary distinction between our work and existing classical simulators, especially for Pauli path simulators, lies in their respective purposes. Specifically, our work focuses on developing a shadow state predictor to predict the mean values of a bounded-gate quantum circuit for new rotation angles and observables. In contrast, existing classical algorithms aim to efficiently simulate the mean values of a bounded-gate quantum circuit for varying rotation angles while keeping the observables fixed. This fundamental difference—prediction versus estimation—necessitates distinct theoretical analyses (i.e., the way of using Pauli transfer matrix (PTM) and the truncation method) and underscores unique benefits and advancements our protocol offered in many practically relevant scenarios.
Different roles of PTM. PTM is a widely used tool in quantum computing, with applications ranging from gate set tomography [127] to quantum learning theory [128] and the analysis of quantum circuit dynamics [76, 65, 129, 130]. In the context of simulating the mean values of bounded-gate quantum circuits, PTM is typically combined with Pauli path simulations to compute the contribution of each path [76, 65]. As explained in SM A.2, given an $N$-qubit quantum circuit $U(\bm{x})$ is composed of $d$ $\operatorname{RZ}$ gates and GdG-d Clifford gates, the expectation value of the given observable $O$ under the PTM representation yields
[…]
where the notation $\Phi_{\bm{\omega}}(\bm{x})$ with $\bm{\omega}\in\{0,\pm 1\}^{d}$ refers to the trigonometric monomial basis defined in Eq. (21) and $\operatorname{Tr}(\rho_{\bm{\omega}}O)$ refers to the expectation value of the purely-Clifford circuit for the path indexed by $\bm{\omega}$. For Pauli path simulators, the expectation value for a portion of $\{\bm{\omega}\}$ should be explicitly computed on the classical hardware.
Unlike Pauli path simulators, our proposal adopts a fundamentally different approach to leveraging PTM. Instead of computing nontrivial $\operatorname{Tr}(\rho_{\bm{\omega}}O)$ like Pauli path simulators, we utilize PTM to derive the explicit formula of the bounded-gate quantum state with respect to the input $\bm{x}$, i.e.,
[…]
where $\Phi_{\bm{\omega}}(\bm{x})$ and $\rho_{\bm{\omega}}$ follow the same definitions in Eq. (22). This explicit form inspires us to design the shadow state predictor $\hat{\sigma}_{n}(\bm{x})$ in Eq. (76). Moreover, in Appendix F.2 (proof of Lemma F.2), we prove that given any new input $\bm{x}^{\prime}\in[-\pi,\pi]^{d}$, the expectation value of the shadow state predictor $\hat{\sigma}_{n}(\bm{x}^{\prime})$ matches the truncated target quantum state $\rho_{\Lambda}(\bm{x}^{\prime})$, i.e., $\mathbb{E}[\hat{\sigma}_{n}(\bm{x}^{\prime})]=\rho_{\Lambda}(\bm{x}^{\prime})\equiv\sum_{\bm{\omega},\|\bm{\omega}\|_{0}\leq\Lambda}\Phi_{\bm{\omega}}(\bm{x}^{\prime})\rho_{\bm{\omega}}$.
The above result sets our approach apart from the use of PTM in Pauli path simulators. Specifically, our work leverages PTM to derive the explicit formula of the concept class with respect to the input $\bm{x}$ and subsequently design the shadow state prediction model $\hat{\sigma}_{n}(\bm{x})$, whereas Pauli path simulators such as LOWESA utilize PTM to compute values for each nontrivial path.
Truncation strategy. The low-weight truncation scheme, as with PTM, is also a widely utilized strategy in quantum computing. However, the approach to truncation and the elements being truncated differ substantially across applications, leading to distinct theoretical results and heuristic performances. For Pauli path simulators, low-weight truncation is applied to the frequencies $\bm{\omega}$ or the weights of Pauli operators to achieve computational efficiency while maintaining a low estimation error. In contrast, in our work, low-weight truncation is applied to the frequencies to ensure computational efficiency while minimizing the prediction error. This fundamental distinction in purpose underscores that prior results on Pauli path simulators offer no direct implications for learning protocols, necessitating an entirely new theoretical framework.
Recall that Theorem 1 in the main text establishes that no efficient learning model can predict all bounded-gate quantum circuits. Consequently, the most critical challenge lies in identifying specific classes of bounded-gate quantum circuits that can be efficiently learned—a question that classical simulators do not need to address. To tackle this, we establish a connection between the gradient norm of bounded-gate circuits and the truncation error as indicated in Lemma F.1. This relation warrants that the proposed learning model can achieve efficient predictions with low prediction error. This novel result, which is absent in Pauli path simulators, highlights the uniqueness of our approach and opens new avenues for developing efficient learning models tailored to predicting the mean values of bounded-gate quantum circuits.
Unique benefits of our learning model in two scenarios. As explained previously, our method focuses on the shadow state prediction, while LOWESA centers on Heisenberg simulations. These differences define the unique strengths of our model in the following two scenarios.
  • (i)
    Circuits with many Clifford gates. Classical Pauli path simulators, such as LOWESA, must explicitly simulate Clifford gates, which, although theoretically efficient ($o(N^{2})$ complexity with $N$ being the qubit count), can be computationally expensive in terms of wall-clock time. In contrast, our learning model is insensitive to the number of Clifford gates, as these gates do not affect the learning process. We validate this advantage through numerical simulations in SM K.5, demonstrating superior performance in this regime.
  • (ii)
    Circuits with unknown system noise. Classical simulators, including both tensor network simulators and Pauli path simulators, rely on assumptions about noise strength when simulating the dynamics of noisy quantum circuits. However, these assumptions generally do not align with the actual noise present in the system. In contrast, a learning-based approach (such as the extension of our model) can automatically capture and adapt to inherent system noise by leveraging the dataset collected from the target quantum system, offering a more flexible and accurate alternative.
SM K More numerical simulations
@@
In this section, we demonstrate more simulation details omitted in the main text. Specifically, in SM K.1, we provide heuristic evidence about the required truncation value $\Lambda$ in practice. In SM K.2, we provide more simulation details about predicting properties of rotational $N$-qubit GHZ states. Then, in SM K.3, we illustrate more simulation results about the global Hamiltonian simulation task. Next, in SM K.4, we demonstrate how to use the proposed ML model to enhance variational quantum algorithms, including variational quantum Eigen-solver and quantum neural networks. Last, in SM K.5, we numerically explore the performance of the proposed ML model over and LOWESA when processing quantum circuits with many Clifford gates.
K.1 Heuristic evidence about the truncation value $\Lambda$
In this subsection, we conduct a systematic analysis to heuristically understand the behavior of $\Lambda$, or equivalently $|\mathfrak{C}(4C/\epsilon)|$. The results provide insights into the conditions under which the proposed ML model can operate efficiently.
The scaling behavior of $\mathfrak{C}(\Lambda)$. We begin by examining how the cardinality of the truncated frequency set $\mathfrak{C}(\Lambda)$ defined in Theorem 2 of the main text scales with the input dimensions $d$ (i.e., the number of $\operatorname{RZ}$ gates) and the truncation thresholds $\Lambda$. Mathematically, we have
$p\equiv|\mathfrak{C}(\Lambda)|=\sum_{k=0}^{\Lambda}\binom{d}{k}2^{k}.$ (144)
Supplementary Fig. 5: The problem-dependent behavior of the truncation threshold $\Lambda$. (a) Scaling behavior of $p$. The x-axis and y-axis represent the truncation thresholds and the cardinality of the truncation set $p\equiv|\mathfrak{C}(\Lambda)|$, respectively. The symbol $\star$’ marks that the cardinality of the truncation set is $p=242$ when $d=5$ and $\Lambda=5$. (b) Quantum circuit for average gradient norm computation. The panel illustrates the quantum circuits used for computing the average gradient norm $C$ when the number of qubits is $N=4$. The dashed box highlights the implementation of hardware-efficient ansatz and the notation ‘$L$’ denotes the number of layers. (c) Scaling behavior of $C$. The simulation results of the average gradient norm $C$ with varied qubit count $N$ and layer number $L$. The vertical bar refers to the variance of the gradient norm.
The scaling behavior is visualized in Fig. 5(a), where $p$ exhibits two distinct phases: in Phase I, for small truncation values $\Lambda$, $p$ scales exponentially with both $\Lambda$ and the input dimension $d$; in Phase II, as $\Lambda$ increases further, the growth of $p$ begins to saturate and eventually converges to a fixed value. Notably, such a scaling imposes a substantial computational burden for large $d$ and $\Lambda$. As shown in the plot, when both $d$ and $\Lambda$ are large, the computational overhead becomes unaffordable, e.g., when $d=50$ and $\Lambda=10$, $p$ is above $10^{12}$.
The relation between $\Lambda$ and the gradient norm $C$. The above analysis indicates that the practical efficiency of the proposed ML model lies in the small-$\Lambda$ regime, especially for the large $d$. According to the relation between $\Lambda$ and the gradient norm $C$ in Theorem 2, this is equivalent to requesting a small value about the ratio $4C/\epsilon$. In this regard, the problem reduces to determining how the average gradient norm $C$ scales in practice.
[…]
and
$U(\bm{x})=\prod_{l=1}^{L}\Big(\bigotimes_{j=1}^{N}\operatorname{RZ}(\bm{x}_{j*(l-1)})\operatorname{RY}(\bm{x}_{j*(l-1)+1})\operatorname{RZ}(\bm{x}_{j*(l-1)+2})\Big)u_{e},$ (146)
where $L$ denotes the number of layers, $u_{e}$ refers to the entanglement layer consisting of $N$ CNOT gates applied sequentially to adjacent qubits, and the dimension of the input yields
d=3NL.d=3*N*L.
The observable interacted with $\rho(\bm{x})$ is fixed to be $O=(X_{1}+Y_{1}+Z_{1})/3$, where $X_{1}$ (or $Y_{1}$, $Z_{1}$) refers to apply the Pauli operator $X$ (or $Y$, $Z$) to the first qubit. An illustration of the employed quantum circuit is depicted in Fig. 5(b).
The hyper-parameter settings are as follows. The number of qubits $N$ varies from 2 to 14, and for each specified qubit count, the number of layers $L$ is varied from 5 to 500. For each configuration, we uniformly and randomly sample the input 500 times to estimate the average gradient norm $C$ and its variance.
The scaling behavior of the average gradient norm is illustrated in Fig. 5(c). An important observation is that, for all layer numbers $L$, the average gradient norm $C$ converges to a small value with the negligible variance as the qubit count $N$ increases. When $N=14$, the value of C with L=50 is around $0.13\pm 10^{-6}$. Although the concrete value of $C$ depends on the employed ansatz $U(\bm{x})$ and the qubit count $N$, the obtained simulation results support a reasonable conjecture that $C\ll 1$ as the number of qubits increases.
The above findings validate the potential of the proposed ML model to predict interested properties of large-qubit circuits. According to the relation $\Lambda=4C/\epsilon$ and the scaling behavior of $\Lambda$ exhibited in Fig. 5, we conclude that the primary factor influencing the applicability of the proposed ML model is the error threshold $\epsilon$. That is, when the given task requires high-precision predictions such that $\Lambda=4C/\epsilon$ yields a very large constant, the proposed ML model will encounter computational issues; otherwise, the proposed ML model can be efficient.
K.2 Numerical simulations of $N$-qubit rotational GHZ states
@@
[…]
The circuit implementation of this class of quantum states is visualized in Fig. 6(a).
Supplementary Fig. 6: More simulation results of rotational N-qubit GHZ states. (a) Circuit implementation of the rotational $N$-qubit GHZ states. The three $\operatorname{RY}$ gates apply to the first qubit, the $(N/2)$-th qubit, and the last qubit. (b) Prediction error. The root mean squared (RMS) error of the trained ML model with varied truncation $\Lambda$, the shot number $T$, and the qubit count NN.
@@
Prediction error versus varied number of qubits $N$. We append more simulation results about how the prediction error of the proposed ML model depends on the shot number $T$, the truncation $\Lambda$, and the qubit counts $N$. To be specific, we fix the number of training examples to be $n=500$ and collect these training examples under different qubit count $N$, where the maximum shot number is set as $T=150$. Fig. 6(b) visualizes the root mean squared (RMS) prediction error under different settings on $10$ test examples. The achieved results indicate that the performance of the proposed ML model is dominated by the truncation number $\Lambda$ (i.e., the dimension of classical inputs $d$) and not sensitive to the number of qubits, which echoes our theoretical analysis. More precisely, when $\Lambda\geq 2$ and $T\geq 100$, the prediction error attains a very low value for both $N=10,30,60$.
K.3 More details of synthetic global Hamiltonian simulation
Here we conduct a much more difficult task compared to the one shown in the main text. In particular, we apply the proposed ML model to predict the magnetization with $\braket{\bar{Z}}$ when $d=30$. The number of training examples and the snapshot for each example are set to $n=50000$ and $T=50$, respectively. In this scenario, full expansion becomes computationally infeasible, where the cardinality of the frequency set is $3^{30}$. The corresponding cardinality of the frequency set is $|\mathfrak{C}(\Lambda=1)|=61$, $\mathfrak{C}|(\Lambda=2)|=1801$, $|\mathfrak{C}(\Lambda=3)|=472761$, and respectively. When evaluating the statistical performance of the proposed ML model with $n<50000$, we sample a subset from the whole dataset using different random seeds. The random seeds for all relevant simulations are set as $12345$, $22222$, $33333$, $44444$, and $55555$, respectively.
RMS prediction error. Fig. 7 demonstrates the RMS prediction error of our proposal with the varied number of training examples, i.e., $n\in\{10,10^{2},10^{3},10^{4}\}$. An immediate observation is that with the increased $n$, the averaged prediction error continuously decreases for all settings of $\Lambda$. In addition, the achieved results hint that the proposed model with a larger truncation value $\Lambda$ may require more training examples to surpass the one with a smaller truncation value. For instance, the proposed ML model with $\Lambda=2$ attains a better performance compared to $\Lambda=1$ when $n=10^{4}$. It is expected that the model with $\Lambda=3,4$ can attain better performance by further increasing $n$.
Supplementary Fig. 7: Prediction error on the state evolved by $60$-qubit global Hamiltonians with $d=30$ with different sizes of the training dataset. The meaning of notations in the box plot is as follows. The label $\Lambda=a$’ stands for that the truncation value is set as $a$. The $x$-axis refers to the varied number of training examples $n$, ranging from $n=10$ to $n=10^{4}$. The $y$-axis refers to the achieved RMS prediction error. All results are obtained using the same random seeds.
For completeness, we also append the corresponding standard error in Table 3. Particularly, for all settings with $\Lambda\in\{1,2,3,4\}$, the averaged RMS prediction error and the corresponding standard error continuously decrease with the increased number of training examples.
Supplementary Table 3: Simulation results of predicting the magnetization of states evolved by the 60-qubit global Hamiltonian.
@@
The role of shot number $T$. Here we conduct numerical simulations to explore how the shot number $T$ influences the prediction error. Specifically, we fix the number of training examples as $n=500$ and the truncation value as $\Lambda=1$, but vary the shot number from $T=50$ to $T=500$. Each setting is repeated five times to collect the statistical results, where the randomness stems from distilling different training datasets from $600$ training examples. The simulation results are summarized in Table. 4, where the averaged prediction error does not decrease with the increased number of measurements. This phenomenon echoes our theoretical analysis, suggesting that once the shot number exceeds a certain threshold, it does not heavily affect the performance of the proposed model.
Supplementary Table 4: Simulation results of predicting the magnetization of states evolved by the 60-qubit global Hamiltonian with varied shot numbers.
@@
$n=500$ & $\Lambda=1$ $0130\pm 0.0109 0.0$$0251\pm 0.0038$
K.4 More simulation results of enhanced variational quantum algorithms by pretraining
@@
Algorithmic implementation. Recall that contemporary quantum devices encounter limitations such as connectivity, gate fidelities, and coherence time. To overcome these constraints, experimentalists typically adopt tailored ansatzes constructed from the native gate set from the specified quantum device to implement various variational quantum algorithms [132, 133]. Essentially, these ansatzes maintain the consistent gate layouts while adjusting parameters for various computational tasks, as delineated by the framework formulated in Eq. (3) of the main text. This intrinsic relation warrants the use of the proposed ML model to enhance plenty of variational quantum algorithms associated with device-specific or task-specific ansatzes.
The learning framework is summarized in Alg. 1. For a given quantum device and a specified ansatz, the learner first constructs the training dataset following the outlined procedure in the main text. Once the dataset is prepared, the proposed ML model serves as a surrogate to optimize variational quantum algorithms in which the optimization correspond to minimize the output value of the surrogate, obviating the need for direct implementation on the specified quantum device. Since the optimization process (i.e., the third step) is conducted entirely on classical processors, it significantly reduces the overhead associated with accessing sparse quantum devices in the contemporary era.
[…]
@@
In what follows, we first briefly explain how our approach supports two crucial classes of variational quantum algorithms: the variational quantum Eigen-solvers (VQEs) and quantum neural networks (QNNs). Then, in SM K.4.1, we present numerical simulations demonstrating the application of the proposed pretraining strategy to enhance a 3-qubit ansatz for both ground-state energy estimation and binary classification tasks. To demonstrate the potential of the proposed pretraining strategy at the use-case level, we employ it to enhance VQEs when processing a class of transversefield Ising models (TFIMs) up to 50 qubits in SM K.4.2. In addition, we conduct numerical simulations to illustrate how the proposed pretraining strategy can advance VQEs when handling a class of $\mathsf{H}_{5}$ molecules in SM K.4.3. Note that the learning scheme presented below can be readily extended to wide applications covered by variational quantum algorithms.
Remark. The focus on TFIMs and $\mathsf{H}_{5}$ molecules stems from the fact that they are widely studied in the literature, garnering attention in quantum many-body physics and quantum chemistry, spanning algorithmic developments to experimental investigations. In particular, TFIM serves as a prototypical model for understanding quantum phase transitions, non-equilibrium dynamics, and critical phenomena in condensed matter systems [134]. Its simplicity, coupled with its rich physics, makes it a cornerstone for testing quantum algorithms [135, 136] and benchmarking quantum simulators [137, 15]. In addition, the hydrogen molecule (e.g., $\mathsf{H}_{2}$ and $\mathsf{H}_{5}$) represents a small yet nontrivial system that captures essential features of electron correlation and quantum entanglement in molecular systems [138]. It is widely used to benchmark quantum computational methods in solving quantum chemistry problems with near-term quantum devices [139, 140].
@@
[…]
@@
[…]